# Two-Time Relativistic Bohmian Model of Quantum Mechanics

### v9

@@ IX Final remarks
The proposed theory arises as a result of taking literally Feynman’s sum over history interpretation. The postulation of a new time parameter is a much simpler expedient, from a mathematical point of view, than the criterion of the Bohmian theories mentioned in the introduction, which predict a convoluted space-time and violate the spirit of Special Relativity. The theory is able to satisfactorily explain the empiricism of Quantum Mechanics, leading interference, collapses and self-interactions (called suspiciously virtualvirtual in the standard view) to a deterministic explanation. Furthermore, TTBM provides testable effects beyond the standard viewpoint and exciting development prospects involve spin and dark matter.
Of course, the theory is in its infancy and we have pointed out some gap that need to be filled. An important issue concerns the integration of TTBM with the Field Theory, a pending task probably capable of clarifying important elements.
@@
$\frac{\nabla^{2}\psi}{\psi}=\frac{\nabla^{2}R}{R}-\frac{p^{2}}{\hbar^{2}}+\frac{{\rm i}}{R^{2}\hbar}\vec{\nabla}\cdot(R^{2}\vec{p}_{c})=0.$ (42)

### v10

@@ IX Final remarks
The proposed theory arises as a result of taking literally Feynman’s sum over history interpretation. The postulation of a new time parameter is a much simpler expedient, from a mathematical point of view, than the criterion of the Bohmian theories mentioned in the introduction, which predict a convoluted space-time and violate the spirit of Special Relativity. The theory is able to satisfactorily explain the general empiricism of Quantum Mechanics, leading interference, collapses and self-interactions (suspiciously called virtualvirtual in the standard view) to a deterministic explanation. Furthermore, TTBM provides testable effects beyond the standard viewpoint and exciting development prospects involve explanation of spin and dark matter.
Of course, the theory is in its infancy and we have pointed out some gaps that need to be filled. The most important seems to be the derivation, through approximations, of the various standard spin relativistic wave equations, in primis the Dirac one.
@@
$\frac{\nabla^{2}\psi}{\psi}=\frac{\nabla^{2}R}{R}-\frac{p^{2}}{\hbar^{2}}+\frac{{\rm i}}{R^{2}\hbar}\vec{\nabla}\cdot(R^{2}\vec{p}_{c}).$ (42)